Find the Tangent and Normal line to the given curve. 9x^3 - y^3 = 1 at (1,2)


1
Expert's answer
2021-01-04T20:39:48-0500

Solution:

Tangent will have a slope of "\tfrac{dy}{dx}"

Normal will have a slope of "- \tfrac{dx}{dy}"


"9x^3-y^3=1"


Let's confirm that it does pass through (1 , 2)

"9*1^3-2^3=1"

9 - 8 = 1

1 = 1

Confirmed

"9x^3-y^3=1"


Derive implicitly

"27x^2dx-3y^2dy=0"

"9x^2dx-y^2dy=0"

"9x^2dx=y^2dy"

"x=1,y=2"


"9*1^2dx=2^2dy"

"9dx=4dy"

"\tfrac{9}{4}=\tfrac{dy}{dx}"


"-\tfrac{4}{9}=-\tfrac{dx}{dy}"


Point-slope form:  y - k = m * (x - h), where m = slope and (h , k) is a point on the line.


Tangent: m = "\tfrac{9}{4}" , (1 , 2)

y - 2 = ("\tfrac{9}{4}" ) * (x - 1)


Normal:  m = -"\tfrac{4}{9}" , (1 , 2)

y - 2 = (-"\tfrac{4}{9}" ) * (x - 1)